OpenStudy (anonymous):

x^2-25/x-5, simplified its x+5, my question are what are the restrictions on it???? I really need help here

OpenStudy (anonymous):

its either 5 or -5

OpenStudy (unklerhaukus):

what would happen if x=5,

OpenStudy (anonymous):

it would be 10

OpenStudy (anonymous):

x2-25 = (x+5) (x-5) ok?

OpenStudy (unklerhaukus):

\[\frac{x^2-25}{x-5}\rightarrow\frac{5^2-25}{5-5}\]

OpenStudy (anonymous):

(x+5)(x-5) / (x-5) = (x+5)

OpenStudy (anonymous):

x cant equal 0 and if x = -5 it would be 0 so the restriction is -5?

OpenStudy (unklerhaukus):

\[\frac{x^2-25}{x-5}\rightarrow\frac{5^2-25}{5-5}\rightarrow\frac 00\]

OpenStudy (cwtan):

anything/0= undefined

OpenStudy (anonymous):

so its 5

OpenStudy (unklerhaukus):

for x=-5\[\frac{x^2-25}{x-5}\rightarrow\frac{(-5)^2-25}{-5-5}\rightarrow\frac{25-25}{-10}=\frac0{-10}=0\]

OpenStudy (anonymous):

so I am confused is it 5 or not?

OpenStudy (anonymous):

here x can not be equal to 5, else it will take indeterminant form (0/0) ... and for rest of the values it simplifies to x+5

OpenStudy (unklerhaukus):

if you try to plot \(y(x)=\frac00\) on a graph it dosent make a lot of sense, you have en-counted a hole in the function